Suppose that market demand is given by the equation qd=111.00−p, and market supply is given by the equation qs=p−15.00. If the government imposes a price ceiling on this good at a price of $25.00, what would be the change in producer's surplus relative to the market equilibrium?


1
Expert's answer
2021-10-06T09:41:06-0400

Solution:

Derive equilibrium price and quantity:

Qd = 111 – P

Qs = P – 15

At equilibrium: Qd = Qs

111 – P = P – 15

111 + 15 = P + P

126 = 2P

P = 63

Qd = 111 – 63 = 48

Equilibrium price = $63

Equilibrium quantity = 48 units


When the price ceiling is set at $25, it means it is below the equilibrium price of $63.

When the price ceiling is set below the equilibrium price, quantity demanded will exceed quantity supplied, resulting in excess demand or shortages in the market.


Therefore, producer surplus relative to the market equilibrium will clearly fall since the price ceiling will transfer some producer surplus to consumers.